### v2

# Unifying Multitrack Music Arrangement via Reconstruction Fine-Tuning and Efficient Tokenization

Abstract
Automatic music arrangement streamlines the creation of musical variants for composers and arrangers, reducing reliance on extensive music expertise. However, existing methods suffer from inefficient tokenization, underutilization of pre-trained music language models (LMs), and suboptimal fidelity and coherence in generated arrangements. This paper introduces an efficient multitrack music tokenizer for unconditional and conditional symbolic music generation, along with a unified sequence-to-sequence reconstruction fine-tuning objective for pre-trained music LMs that balances task-specific needs with coherence constraints. Our approach achieves state-of-the-art results on band arrangement, piano reduction, and drum arrangement, surpassing task-specific models in both objective metrics and perceptual quality. Additionally, we demonstrate that generative pretraining significantly contributes to the performance across these arrangement tasks, especially when handling long segments with complex alignment. [^1] Demos available at ae4rtjsyr.github.io. All code will be released upon acceptance.
1 Introduction
Music arrangement, the art of adapting musical compositions for different performance contexts, plays a fundamental role in music creation by enabling artistic reinterpretation and expanding the accessibility of musical works. It is ubiquitous across various musical domains, from professional music production and education to live performance and digital content creation. However, manual arrangement requires extensive expert knowledge of harmony, instrumentation, and style-specific conventions, making it a highly specialized and demanding process. Automating this task could make arrangement more accessible to individuals with limited musical training and allow composers to focus more on the creative aspects by relieving them of technical groundwork.
Despite the potential benefits, existing methods for automatic arrangement face significant challenges, leading to substantial quality gaps that hinder practical applications. One critical issue lies in tokenization, the foundation of any unconditional or conditional generative music modeling. The commonly adopted REMI tokenizer Huang and Yang (2020) often results in excessively long sequence lengths, a known bottleneck for symbolic music generation Hsiao et al. (2021). Moreover, as an REMI’s extension to multitrack music, REMI+ von Rütte et al. (2023) contains redundant information and suboptimal design choices that lead to the content fragmentation issue in the tokenized sequences, burdening both generative modeling and arrangement performance.
Another limitation of prior arrangement methods is their reliance on task-specific model structures. While generative pre-training has demonstrated great potential in related fields such as NLP, integrating similar approaches into these highly specialized structures is not straightforward. Consequently, these methods fail to fully leverage the generative power of pre-trained music language models (LMs). Given the scarcity of annotated music data, this limitation further constrains their effectiveness.
Finally, existing models often struggle with fidelity and coherence. Specialized arrangement models Zhao et al. (2023, 2024) frequently fail to preserve melody and texture, retaining only harmony, and thus are perceived by users as lacking resemblance to the original piece. Also, our experiments show that even the state-of-the-art piano reduction model Terao et al. (2023) exhibit suboptimal arrangement coherence since different musical segments are handled independently.
To address these limitations, we propose a novel framework for automatic symbolic music arrangement. First, we design a more efficient tokenizer to solve the content fragmentation issue and reduce sequence length and redundancy. Second, we explore the potential of sequence-to-sequence fine-tuning with our pre-trained music LM, enabling task specific models to harness the power of generative pre-training. This approach allows flexible adaptation to various arrangement tasks with limited annotated data. Finally, we validate our unified methodology across multiple tasks, including band arrangement, piano reduction, and drum arrangement.
Our key contributions are summarized as follows:
  • We propose an efficient tokenizer for multitrack music that reduces sequence length and redundant information, improves effectiveness for unconditional pre-training, enhances arrangement performance, and potentially extensible to general music generation tasks.
  • We introduce a self-supervised reconstruction objective to effectively adapt pre-trained music LMs to various music arrangement tasks, ensuring both task-specific requirements and coherence.
  • We achieve state-of-the-art performance on band arrangement, piano reduction, and drum arrangement tasks, surpassing task-specific models in both objective and subjective evaluations.
  • Through ablation studies, we demonstrate how generative pre-training significantly improves bar-to-bar transformations and plays a crucial role in processing longer segments, where models without pre-training fail.
2 Related Work
2.1 Automatic Arrangement in Multitrack Symbolic Music
Music arrangement encompasses various tasks like chord progression creation Yi et al. (2022), lead sheet orchestration Wang et al. (2024), and composition adaptation Zhao et al. (2024). This paper focuses on adapting existing music to new instruments or ensembles, which is crucial for multi-style music generation and style transfer.
Earlier supervised approaches trained end-to-end models using parallel data (e.g., piano-to-orchestra Crestel and Esling (2016) or band-to-piano Terao et al. (2022, 2023)), which are extremely challenging to prepare and inherently restrict the direction of arrangement. Classification-based methods such as Dong et al. (2021) attempt to predict instrument labels for each note, but they lack creativity by generating no new content and suffer from inflexibility, being incapable of handling unseen instrument combinations. Recent self-supervised methods like Q&A Zhao et al. (2023, 2024) offer more flexible solutions without requiring parallel datasets or fixed instrument set assumptions. However, Zhao et al. (2023) assumes prior knowledge of track-wise output distributions (i.e., note density and pitch histograms), which is impractical for arrangement tasks since such knowledge is unavailable before generation. While Zhao et al. (2024) attempts to solve this by introducing a separate prior model to estimate the track-wise distributions, this approach comes at the cost of compromised arrangement fidelity. Additionally, infilling-based models such as Composer’s Assistant Malandro (2023, 2024), although capable of handling additive arrangement scenarios such as drum arrangement, do not preserve the contents of the original composition in the generation output and are thus fundamentally unsuited for band arrangement, where the primary aim is to reinterpret existing musical content.
2.2 Generative Symbolic Music Models
Transformer networks have advanced music generation tasks like chord-to-melody generation Madaghiele et al. (2021), accompaniment arrangement Ren et al. (2020), and music inpainting Malandro (2023, 2024). However, these models remain task-specific, with limited exploration of unconditional pre-training and task-specific fine-tuning. Regarding controlled generation, existing works primarily focus on using simple high-level attributes (style Choi et al. (2020); Lu et al. (2023), structure Zhang et al. (2022), and sentiment Makris et al. (2021)), with minimal investigation of music-based content conditioning. As large symbolic music LMs emerge Dong et al. (2023); Qu et al. (2024); Wu et al. (2024), it becomes crucial to explore more complex control signals, such as music itself, and expand their application to broader scenarios.
2.3 Symbolic Music Tokenizers for Transformers
Transformer-based music modeling requires converting data into token sequences, which is challenging due to music’s multi-stream nature. Many tokenizers treat music as a single stream using MIDI-like note-event encoding, representing each musical note with multiple tokens (onset time, pitch, duration, velocity, instrument). Some use absolute timing Huang et al. (2018); Gardner et al. (2021); Zeng et al. (2021); Ens and Pasquier (2020), while others use metrical durations Ren et al. (2020); Huang and Yang (2020); von Rütte et al. (2023). Among them, the REMI Huang and Yang (2020) tokenizer is widely adopted for single-track music and has been extended to multitrack in von Rütte et al. (2023) by adding an instrument token as an attribute of each note.
However, such extension, along with all approaches that strictly order notes temporally for multitrack music, has an inherent limitation—content fragmentation, i.e., notes from one instrument being interleaved with notes from other instruments playing simultaneously. As shown in Figure 1(a), the content of instrument i-29 (distorted electric guitar, highlighted in orange) spreads across the entire sequence, interleaved by notes from instrument i-80 (synth lead, highlighted in red). Consequently, tokenization results can be highly inconsistent for same instrument-wise contents once its context (interleaving contents) differs, bringing extra difficulty for generative modeling.
2.4 Prompt-based Fine-Tuning for Conditional Generation
Prompt-based fine-tuning in NLP Liu et al. (2023) uses control tokens to guide pre-trained LM outputs, enabling attribute control over style Sennrich et al. (2015), length Kikuchi et al. (2016), and pronunciation Ou et al. (2023). Combining such strategy with LMs pre-trained with standard left-to-right language modeling objective Radford et al. (2019); Brown et al. (2020) can effectively generates fluent, condition-aligned outputs, providing insights for the music arrangement problems, where similar pre-training and control mechanisms can potentially generate music that is musically appealing and meets task-specific requirements.
3 Method
To bridge the gaps in previous research, we adopt a pre-train and fine-tune paradigm for the arrangement task, design an efficient multitrack music representation for Transformer networks that makes both pre-training and fine-tuning more effective, and develop a unified objective for music-to-music transformation that addresses both task-specific requirements and coherence. We test the effectiveness of this approach on multiple typical music arrangement tasks.
(a) REMI+ tokenization, demonstrated with REMI-z vocabulary.
(b) A REMI-z bar sequence containing four track sequences.
Figure 1: REMI-z and REMI+ tokenization for a same bar. Content of the same instruments are highlighted with same color.
(a) Detailed settings for each arrangement task. The symbol $\bigoplus$ denotes concatenation of component sequences.
(b) An example of the tokenized sequence for a band arrangement task. Special tokens $<$$\tt{SEP}$$>$, $<$$\tt{INSTRUMENT}$$>$, $<$$\tt{CONTENT}$$>$, and $<$$\tt{HISTORY}$$>$ are used to separate different components. Notes of each instrument are highlighted with different colors.
Figure 2: Music segments are decomposed into three subsequences: instruments, content, and target-side history. These components form the condition sequence, with the original music as the target sequence. The model is trained to reconstruct the music from these components.
3.1 REMI-z Representation
Meaning Token X’s range
Instrument type i-X 0$\sim$128
Note’s within-bar position o-X 0$\sim$127
Note’s pitch (non-drum) p-X 0$\sim$127
Note’s pitch (drum) p-X 128$\sim$255
Note’s duration d-X 0$\sim$127
End of a bar b-1 -
Time signature s-X 0$\sim$253
Tempo t-X 0$\sim$48
Table 1: REMI-z vocabulary with pitch token distinctions
Instruments have inherent syntactic rules governed by their physical properties and conventional playing patterns. A good arrangement, or in broader terms, any musical composition, must ensure each instrumental part is both executable and idiomatic. However, as mentioned earlier, time-ordered tokenizers struggle with instrument-wise content fragmentation, making it difficult to model these syntactic rules. To address this issue, we propose REMI-z, a tokenization scheme that does not sort notes strictly by time but instead prioritizes the sequence continuity of content for each instrument.
The vocabulary of REMI-z is shown in Table 1, similar to that of REMI+ von Rütte et al. (2023) (detailed in Appendix A). The key innovation lies in how notes are organized. REMI-z tokenize arbitrary-length MIDI files into a list of bar sequences, each ending with a special end-of-bar token (b-1). Within each bar, notes are further grouped into track sequences, where each track corresponds to a specific instrument. Each track sequence begins with an instrument token specifying the instrument type, followed by all notes played by that instrument within the bar. Track sequences are ordered by their average pitch (from high to low), and within each track sequence, notes are represented as triplets of (position, pitch, duration). These notes are sorted first by position (left to right) and then by pitch (high to low) before being flattened into a 1-D sequence. If two adjacent notes share the same position, the second position token is omitted.
The primary difference between REMI-z and REMI+ is that in REMI-z, notes belonging to the same instrument are grouped together within each bar (as in Figure 1(b)), significantly alleviating the content fragmentation issue and facilitating instrument-specific syntactic modeling. Another advantage is that the sequence length is reduced, as instrument tokens appear only once per track sequence in REMI-z, whereas REMI+ needs an instrument token for each note.
We adopt REMI-z tokenization to pre-train a Transformer decoder using a standard next-token-prediction objective on a large unlabeled corpus and then proceed to convert the pre-trained model to task-specific models with the objective in the following section.
3.2 Arrangement Fine-Tuning
Music arrangement involves adapting given musical content to a new set of instrumental constraints. With a well-trained content- and instrument-controllable generation model, novel arrangements can be generated from the original composition by assigning new instrumentation constraints at inference time. To train this capability without relying on parallel data, such as multiple versions of the same song arranged for different instrumentations, we propose a context-aware self-supervised segment-level reconstruction objective. Let $y^{(t)}$ denote the $t$-th segment of a music piece. The fine-tuning objective is defined as:
$\mathcal{L}(\theta)=-\log\,p_{\theta}\big{(}$ $\mathcal{T}_{\text{task}}(y^{(t)})\,\big{|}\,\mathrm{I}(\mathcal{T}_{\text{task}}(y^{(t)})),$
$\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)})),\,\mathcal{T}_{\text{task}}(y^{(t-1)})\big{)},$ (1)
where $\theta$ represents the model parameters, $\mathrm{I}(\cdot)$ and $\mathrm{C}(\cdot)$ extract instrument and content information, $\mathcal{S}_{\text{task}}$ and $\mathcal{T}_{\text{task}}$ represent selecting task-specific source and target tracks, and $y^{(t-1)}$ provides target-side history. Equation 3.2 is implemented with a standard LM objective to the sequence [condition]<sep>[target], as illustrated in Figure 2(b), with cross-entropy loss computed only on the target subsequence.
During training, the instrument, content, and history conditions are derived from the same music being reconstructed. The model is trained to reconstruct the original music, learning to interpret the given musical content with a specific instrument set in a certain context. During inference, users can flexibly define instrument constraints, while the content comes from the music to be arranged, and history from the model’s previously generated outputs.
The instrument condition, derived from $\mathrm{I}(\cdot)$, specifies the desired instruments to be used in the segment. During training, it includes all instrument tokens present in the target-side sequence. Additionally, the relative positions of tokens within this sequence encode important information about the voice relationships between instruments: tokens earlier in the sequence correspond to instruments with higher average pitches, and vice versa. At inference, users have the flexibility to specify both the instruments and their voice relationships for each segment.
The content condition from $\mathrm{C}(\cdot)$ represents the musical notes from the original composition, encoded as a content sequence. This sequence provides precise information about what notes are played and their timing, but without specifying instruments. To construct the content sequence, the following steps are applied: (1) remove all instrument tokens from REMI-z, (2) sort notes strictly by their temporal positions, (3) sort notes within the same position by pitch in descending order, and (4) merge duplicate notes with identical positions and pitches. This representation ensures a clean, time-ordered view of the musical content of a multitrack segment without any instrument-related information.
Lastly, we introduce a history condition $\mathcal{T}_{\text{task}}(y^{(t-1)})$ that encourages contextual coherence between segments during the arrangement process to tackle the style inconsistency issue brought by segment-level processing, inspired by the success of historical context in machine translation Tiedemann and Scherrer (2017). During training, the history is provided through teacher-forcing using the ground truth REMI-z sequence from the previous segment’s target sequence, while at inference time, it is generated autoregressively based on the model’s output from the preceding segment. This mechanism ensures that each segment is generated with awareness of its musical context, leading to coherent arrangements.
3.3 Tasks
We evaluate our method on the following tasks, which are representative of typical arrangement scenarios in music composition studies. Each task assesses different aspects of model capabilities.
3.3.1 Band Arrangement
This task involves arranging an existing piece of music for arbitrary combinations of instruments. The model must understand the properties and typical playing styles of each instrument to allocate or generate new notes appropriately. In this setup, both $\operatorname{\mathcal{S}_{\text{task}}}$ and $\operatorname{\mathcal{T}_{\text{task}}}$ are defined as identity transformations, directly returning the original REMI-z sequence, i.e., $\operatorname{\mathcal{S}_{\text{task}}}(y)=\operatorname{\mathcal{T}_{\text{task}}}(y)=y$. Note that we focuses exclusively on pitched instruments and pre-remove drum tracks from $y$ due to REMI-z’s drum-specific pitch tokens. Drum arrangement is addressed as a separate task.
To foster creativity, we randomly delete a small proportion of tracks in $\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)}))$ during training, encouraging the model to generate new notes that are compatible with the musical context but absent from the input music during inference. Special care is taken to preserve the melody during random deletions by ensuring that the track with the highest average pitch remains untouched. Additionally, duration tokens are removed from the content sequence to allow the model to generate durations that best suit the playing styles of the desired instruments, rather than merely copying from inputs. The segment length is set to 1 bar.
3.3.2 Piano Reduction
This task focuses on simplifying multi-instrumental musical pieces into solo piano arrangements that ensure pianistic playability while preserving the original musical essence, i.e., harmonies and textures. In the training process, $\operatorname{\mathcal{T}_{\text{task}}}$ is defined to extract the original piano tracks, while $\operatorname{\mathcal{S}_{\text{task}}}$ is an identity transformation. To ensure the piano part is sufficiently prominent in the training data, we filter the dataset to include only segments where the piano pitch range constitutes more than 40% of the total pitch range, reducing the dataset to approximately 40% of its original size. The segment length is set to 1 bar. Additionally, drum tracks are also pre-removed from $y$.
3.3.3 Drum Arrangement
This task involves creating a drum track for songs that lack one. The model needs to recognize the groove of the music and enhance it using the drum set, and further, handle transitions between musical phrases to drive the music forward and make it more engaging, which consequently requires a better understanding of musical structure. Here, $\operatorname{\mathcal{S}_{\text{task}}}$ extracts tracks belonging to pitched instruments, while $\operatorname{\mathcal{T}_{\text{task}}}$ extracts the drum track. A longer segment length of 4 bars is used for this task, as drum patterns generally span across multiple bars.
4 Experiments
4.1 Implementation Details
Our model, an 80M-parameter decoder-only Transformer, has a hidden dimension of 768, 12 layers, 16-head attention, and a context length of 2048 tokens (around 8$\times$ the longest bar in our dataset). The model first undergoes a standard LM pre-training, and then was fine-tuned with the proposed objective. Pre-training used four RTX A5000 GPUs (batch size 12, 1 epoch), while fine-tuning used a single A40 GPU (variable batch size, 3 epochs). Training was done with two publicly accessible datasets. Pre-training adopted the Los-Angeles-Project dataset Lev (2024) (405K MIDI files, 4.3B tokens after REMI-z tokenization, 2% validation split) and fine-tuning was done with Slakh2100 Manilow et al. (2019) (1,289 training, 270 validation, 151 test MIDI files), featuring 34 pitched instruments and drums, with $\geq$5 tracks per piece. Detailed hyperparameter settings are in Appendix B.
4.2 Tokenizer Efficiency Comparison
We compare REMI+ and REMI-z through theoretical and empirical analyses. First, we tokenize the Slakh2100 dataset using both methods and evaluate compactness via the average number of tokens per bar ($\bar{T}_{\text{bar}}$) and per note ($\bar{T}_{\text{note}}$). Lower values indicate reduced computational costs during training and autoregressive decoding.
Next, we quantify token sequence complexity using bar-level Shannon entropy. For a discrete random variable $X$ representing token distributions, entropy of a token sequence $H(X)$ is defined as:
[…]
where $\{x_{1},\dots,x_{N}\}$ are all tokens appeared in the sequence, and $\{P(x_{1}),\dots,P(x_{N})\}$ are their probabilities. We calculate the entropy for each bar sequence to get $H_{\text{bar}}$, and then average across all bars in the dataset to obtain $\bar{H}_{\text{bar}}$. Lower $\bar{H}_{\text{bar}}$ value indicates less inherent complexity of the bar sequence, and by extension, the multitrack music, reflecting efficient encoding with less redundant information.
Then, we train unconditional LMs on the Slakh2100 dataset and observe the note-level perplexity:
$\text{PPL}_{\text{note}}=\exp\left(-\frac{1}{M}\sum_{j=1}^{M}\log P(n_{j}\mid n_{1:j-1})\right),$ (3)
where $M$ is the number of notes in a bar, and $P(n_{j}\mid n_{1:j-1})$ aggregates probabilities for all tokens corresponding to note $n_{j}$ (instrument, position, pitch, duration). Unlike token-level perplexity ($\text{PPL}_{\text{token}}$), this metric normalizes by the number of notes, ensuring fair comparison across tokenization schemes. Lower note-level perplexity indicates better modeling of musical notes, as it reflects reduced uncertainty in predicting all associated attributes.
Finally, we assess task-specific performance with different tokenizers, discussed in the next section.
4.3 Task-Specific Evaluation
This section introduces baseline models and metrics used for evaluating the music arrangement models.
4.3.1 Baseline Models
For each task, we compare our model against a state-of-the-art (SOTA) task-specific baseline. For band arrangement, we adopt Transformer-VAE from Zhao et al. (2024), the strongest previously reported model for multitrack arrangement without assumptions on track type or number. It combines Transformer-based long-term and inter-track modeling with a VQ-VAE generation module. For piano reduction, we compare with Terao et al. (2023) (UNet), the most recent work in this area. For drum arrangement, existing drum-specific models (e.g., Barnabò et al. (2023), Dahale (2022)) are unsuitable as they generate drums from a single melody or instrumental track. Instead, we adopt Composer’s Assistant 2 (CA v2) Malandro (2024), a strong track infilling model capable of handling multitrack inputs and generating drum outputs. Baseline’s implementation details are in Appendix B.
To demonstrate the impact of generative pre-training, an ablation variant of our model without pre-training (w/o PT) was used as a baseline. Additionally, we include rule-based baselines to showcase the difficulty of the tasks. For band arrangement, Rule-Based distributes notes evenly by pitch across instruments. For piano reduction, we use Rule-F (flattened multitrack where the piano plays all notes) and Rule-O (original piano track). Note that Rule-O is not considered ground truth since the piano track is likely to not fulfil the $\geq 40\%$ pitch range requirements. For drum arrangement, the original drum track (Ground Truth) is included anonymously for human evaluation as a golden standard.
4.3.2 Segment-level Objective Evaluation
Objective metrics measure similarity between model outputs and target sequences, assuming closer resemblance indicates higher naturalness and musicality. Given the extreme sparsity of note events in the track-wise piano roll (e.g., only 0.04% non-zero elements in the Slakh2100 dataset under 48th-note quantization), F1-based metrics are more suitable than accuracy-based metrics for similarity evaluation. Following Malandro (2024); Terao et al. (2023), we use note-level F1 to measure similarity between model outputs and target sequences. Specifically, we compute Note F1 (correct onset and pitch) and Notei F1 (additional correct instrument prediction), both under 16th-note quantization for fair comparison.
For piano reduction and drum arrangement, the same models are used in objective and subjective evaluations. For band arrangement, models are separately trained without random track deletion to ensure deterministic outputs. Baseline models are also modified by excluding the prior model to remove long-term context hints, ensuring evaluation fairness.
In addition to the modifications described above, we introduce three additional metrics specific to the band arrangement task. First, Instrument Intersection over Union (I-IoU) evaluates the accuracy of instrument control. Second, Voice Word Error Rate (V-WER) measures the similarity in voice features between the generated output and the reference, indicating the effectiveness of the voice control. Third, Melody F1 (Mel F1) calculates Note F1 on the tracks with the highest average pitch in the output and target music, assessing the preservation of melody, which is critical for arrangement fidelity. The detailed computation of these metrics is provided in Appendix C.
Beyond comparing with the SOTA model, the objective evaluation for band arrangement serves two additional purposes: (1) to assess the task-specific performance of different tokenizers and (2) to validate the design of the fine-tuning objective. The band arrangement task is particularly challenging because it imposes no assumptions on the types or numbers of target instruments, making it more representative of the model’s overall capabilities compared to the other tasks.
4.3.3 Song-Level Subjective Evaluation
To complement similarity metrics and account for perceptual quality and creativity, we conducted human evaluations. Full-piece arrangements were generated by all models and evaluated phrase-by-phrase on a 5-point scale (1: very low, 5: very high). For band arrangement, models were tested across three instrument combinations with different complexity: string trio (3 tracks), rock band (4 tracks), and jazz band (7 tracks). Three metrics were used across band, piano, and drum arrangement tasks: Coherence, which evaluates the natural flow of the arrangement and the consistency of each instrument’s performance and style throughout the piece; Creativity, which assesses the degree of innovation in the arrangement under the constraints of the music’s content and style; and Musicality, which measures the overall musical appeal and aesthetic quality of the arrangement. Further details on the metrics, ensemble settings, questionnaire, and evaluation process are in Appendix D.
Task-specific metrics were also introduced. For band arrangement, Faithfulness measures resemblance to the original in melody and overall feel, while Instrumentation assesses the appropriateness of each instrument’s role within the band and their harmony. For piano reduction, Faithfulness is also adopted but without melody preservation requirements, and Playability assesses the feasibility of the generated contents played by human pianists. For drum arrangement, Compatibility measures how well the drum track blends with other instruments, and Phrase Transition assesses the smoothness of transitions between phrases.
Additionally, significance tests were conducted between our model and the SOTA baselines using within-subject (repeated-measures) ANOVA Scheffe (1999). The significance level are showed on the result tables: * for p-values less than 0.05, $\dagger$ for p-values less than 0.01, and $\ddagger$ for p-values less than 0.001.
5 Results
5.1 Effectiveness of REMI-z
Tokenizer $\bar{T}_{\text{bar}}$ $\bar{T}_{\text{note}}$ $\bar{H}_{\text{bar}}$ $\text{PPL}_{\text{note}}$ $\text{PPL}_{\text{token}}$
REMI+ 225.91 4.03 41.68 116.20 3.00
REMI-z (Ours) 151.68 2.77 29.43 84.11 4.50
Table 2: Tokenizer efficiency comparison.
As shown in Table 2, our proposed tokenization method significantly reduces the average sequence length from 225.91 tokens (REMI+) to 151.68 tokens per bar, achieving a 32.9% reduction. The number of tokens per note also decreases from 4.03 (REMI+) to 2.77 with REMI-z. These results demonstrate that our tokenizer generates a more compact representation for multitrack music, which not only reduces the computational complexity for generative modeling but also accelerates autoregressive decoding and potentially facilitates the modeling of long-term dependencies.
Moreover, this increased compactness does not introduce additional complexity. In terms of bar-level Shannon entropy, REMI-z achieves 29.43 bits/token compared to REMI+’s 41.68 bits/token, representing a 29.4% relative reduction. This indicates that sequences tokenized with REMI-z contain less redundant information and can be further encoded with a shorter expected code length, requiring simpler encoding rules for lossless compression and potentially enabling more effective generative modeling.
We confirm this advantage with empirical analysis. The LM trained with REMI-z achieves lower note-level perplexity (84.11 vs 116.20), demonstrating that modeling the basic unit of music—the note—is more effective with REMI-z. This improvement in note-level modeling naturally extends to better performance in generating bars and complete compositions. Conversely, while individual REMI+ tokens are easire to model (-1.5 token-level perplexity), this does not translate to better note-level and song-level performance.
5.2 Band Arrangement
Model I-IOU V-WER Note F1 Notei F1 Mel F1
Transformer-VAE 97.5 35.0 49.5 40.0 24.7
Transformer w/ REMI+ 95.0 18.2 94.4 76.0 68.8
Transformer w/ REMI-z 99.5 9.9 97.8 77.5 77.8
$+$ Pre-training (Ours) 99.8 7.6 97.5 87.0 84.5
-- voice 99.6 17.6 97.2 84.3 81.5
-- history 100.0 9.0 97.6 77.4 79.4
Table 3: Objective evaluation results of the band arrangement task.
5.2.1 Objective Evaluation
We then evaluate REMI-z’s task-specific performance. As shown in Table 3, REMI-z consistently outperforms REMI+ on all metrics. It achieves better instrument control (-4.5% I-IOU), reducing undesired or missing instruments, and more accurate voice control with lower V-WER (-8.3%). This indicates REMI-z enable the model to better follows user-specified instrument constraints. The higher Note F1 and Notei F1 scores demonstrate improved reconstruction quality, highlighting REMI-z’s effectiveness in the arrangement training. The significantly higher Mel F1 (+9.0%) also indicates better melody preservation and arrangement fidelity.
Pre-training further enhances our model’s performance in three major aspects: (1) improved voice control (-2.3% V-WER), indicating better understanding of voice requirement in the instrumental prompts, (2) more accurate per-instrument reconstruction (+9.5% Notei F1), suggesting enhanced comprehension of instrumental roles in ensembles, and (3) better melody preservation (+6.7% Mel F1), leading to improved arrangement fidelity. For detailed knowledge probing analysis of how pre-training aligns learned representations with musical properties, please refer to Appendix E.
Our model consistently outperforms Transformer-VAE across all metrics. The gap is particularly significant in Note F1 (97.5% vs 49.5%) and Notei F1 (87.0% vs 40.0%), indicating Transformer-VAE’s limitation in inferring instrument-specific content within musical contexts. Our model also achieves substantially better melody preservation (+59.8% Mel F1), ensuring higher-fidelity song-level arrangements.
Ablation studies validate our design choices for voice and history conditioning. Excluding relative voice information from the instrument condition sequence leads to degraded voice relationships (-10.0% V-WER) and consequently lower Notei F1 (-2.7%) and Mel F1 (-3.0%). This demonstrates that voice conditioning enables effective voice control and facilitates instrument role inference. Similarly, excluding history from the condition sequence noticeably reduces Notei F1 (-9.6%) and Mel F1 (-5.1%), laying the foundation for generating coherent outputs in song-level inference.
5.2.2 Subjective Evaluation
Model Fa. Co. In. Cr. Mu.
Rule-Based 3.46 3.05 2.89 3.00 3.07
Transformer-VAE 2.65 2.70 2.72 3.00 2.72
Ours 3.77 3.47 3.49 3.40 3.47
w/o PT 3.19 2.82 2.86 2.93 2.75
Table 4: Band arrangement subjective evaluation results. Fa., Co., In., Cr., and Mu. represent Faithfulness, Coherence, Instrumentation, Creativity, and Musicality, respectively.
As shown in Table 4, our model consistently achieves the highest scores across all metrics, demonstrating its capability to generate high-quality arrangements that adhere to the original musical essence while achieving good Coherence, Creativity, and Musicality. Transformer-VAE scores lower than the Rule-Based model, particularly in Faithfulness and Musicality, indicating its limitations in preserving musical essence and expression. This also suggests that the rule-based method can serve as competitive baselines for band arrangement research. Notably, removing pre-training degrades all metrics, highlighting its importance for generation quality.
5.3 Piano Reduction
Model F1 Fa. Co. Pl. Cr. Mu.
Rule-F - 3.93 3.59 3.14 2.96 3.34
Rule-O - 2.75 3.49 4.07 2.62 2.96
UNet 58.3 2.97 2.90 3.47 2.82 2.78
Ours 85.5 3.63 3.64 3.86 3.14 3.48
w/o PT 78.4 2.25 2.58 3.29 2.67 2.26
Table 5: Piano reduction results. The Pl. represents Playability score.
Table 5 presents the piano reduction results. Our model outperforms all baselines in F1 score, Coherence, and Musicality. Unlike Rule-O and Rule-F which excel only at Playability or Faithfulness respectively, our model achieves a balance between these two aspects. Pre-training also proves essential: without it, the model has lower F1 and underperforms UNet from Terao et al. (2023) across all subjective metrics. In contrast, the model with pre-training significantly surpasses UNet in subjective evaluations. Interestingly, F1 scores do not strictly correlate with perceptual quality, as evidenced by the w/o PT model’s higher F1 but lower subjective evaluation scores compared to UNet.
5.4 Drum Arrangement
Model F1 Comp. Co. Tr. Cr. Mu.
Ground Truth 100.0 4.31 4.18 3.36 3.16 3.78
CA v2 20.3 3.82 4.05 2.86 2.58 3.19
Ours 79.3 3.91 4.03 3.77 3.27 3.57
w/o PT 1.2 2.49 2.19 2.21 2.82 2.05
Table 6: Drum arrangement results. Comp. and Tr. represent Compatibility and Phrase Transition score respectively.
As in Table 6, our model significantly outperforms baseline CA v2 in F1 score. On subjective metrics, our model performs comparably with CA v2 on Coherence, and better at all remaining metrics. Particularly, we notice that CA v2 tends to generate repetitive content throughout the song. Conversely, our model excels in handling phrase transitions and demonstrating creativity, achieving a Musicality score comparable to the ground truth. This indicates that our model effectively captures high-level musical structures during generation, introducing variations between conventional drum patterns and producing more engaging arrangements. The model without pre-training fails to perform meaningfully (1.2% F1), with a notably larger performance gap compared to other arrangement tasks. This demonstrates pre-training’s crucial role in sequence-to-sequence tasks when handling longer musical structures, where complex alignments between source and target sequences require enhanced structural understanding.
6 Conclusion
In this paper, we proposed a unified framework for automatic music arrangement that addresses challenges in tokenization, pre-trained model utilization, and arrangement quality. Our tokenizer demonstrates superior efficiency and enhances modeling effectiveness across various generation tasks, while our training objective shows compatibility with diverse music arrangement tasks. Further, LM pre-training significantly enhances downstream performance, particularly for tasks requiring complex sequence alignment. Taken together, our method achieves state-of-the-art results on band arrangement, piano reduction, and drum arrangement tasks, generating natural, faithful, coherent, creative, and musically appealing outputs. The performance advantage of our framework suggests broader applications in music-to-music transformation tasks like infilling and call-and-response generation. Furthermore, our tokenizer’s compact and efficient design makes it a standalone contribution applicable to general multitrack symbolic music generation tasks beyond the scope of arrangement.
Appendix A REMI-z Tokenization
A.1 Vocabulary
Table 1 presents the complete vocabulary of our proposed REMI-z tokenizer, detailing the value ranges for each token type. For instrument tokens, the values correspond to MIDI program numbers, with 128 specifically designated for the drum set. For pitched instruments, pitch token values directly map to MIDI pitch numbers, while drum set pitches are encoded as MIDI pitch + 128. Both position and duration tokens are quantized in units of a 48th note (one-third of a sixteenth note). Time signature and tempo tokens undergo specific quantization before token mapping; for detailed quantization rules, please refer to our code implementation. All tokens listed in the table were utilized during pre-training, with example REMI-z sequences illustrated in Figure 3(b).
During fine-tuning, we simplified the token set by excluding time signature and tempo tokens. This decision was supported by our analysis of the fine-tuning dataset (Slakh2100 Manilow et al. [2019]), where 94.79% of songs use 4/4 time signatures. Consequently, we restricted our fine-tuning to 4/4 songs and omitted time signature tokens, following practices in related works Wang et al. [2022]; Zhao et al. [2024]. We also excluded tempo tokens since tempo adjustments in digital audio workstations are typically handled as a global parameter, affecting only tempo tokens without altering the any other tokens. This simplification assumes that compositional and performance styles remain consistent across different tempos. However, this assumption may not hold for datasets with significant tempo variations. Therefore, for future research requiring time-signature- or tempo-specific characteristics, we recommend including these tokens during fine-tuning.
Velocity tokens were excluded from both pre-training and fine-tuning phases. This decision reflects our focus on compositional quality rather than performance naturalism, aligning with previous approaches in band and piano arrangement studies Zhao et al. [2024]; Terao et al. [2023]. However, if velocity information is deemed crucial for generation, our model architecture readily accommodates the addition of velocity tokens to each note without introducing content fragmentation issues, if the REMI-z note organization order is maintained.
A.2 Example
(a) A 4-bar musical segment with orange arrows illustrating the ‘zig-zag’ encoding order of notes in REMI-z tokenization.
(b) The REMI-z sequence tokenized from the musical segment above. Track sequences are color-coded by instrument: synth lead (red), clean electric guitar (blue), distorted electric guitar (orange), and electric bass (purple).
Figure 3: An example of REMI-z tokenization.
Figure 4: The content sequence obtained by applying the operator $\mathrm{C}(\cdot)$ to the REMI-z sequence shown in Figure 3(b).
We name our tokenization scheme REMI-z for its distinctive “zig-zag” encoding pattern for musical notes within each bar, as illustrated in Figure 3(a). In this scheme, notes are encoded hierarchically: first grouped by tracks (instruments), then organized bar by bar. This track-first approach ensures notes from the same instrument remain clustered together, thereby enhancing the model’s ability to learn instrument-specific patterns. The resulting REMI-z sequence is demonstrated in Figure 3(b). In contrast, REMI+ von Rütte et al. [2023] employs a column-wise encoding strategy, strictly ordering notes by their temporal positions. While this approach effectively captures global temporal relationships, it disperses notes from the same instrument throughout the sequence, potentially complicating instrument-specific pattern learning. We leverage this temporal ordering advantage of REMI+ in our content sequence (excluding instrument tokens), as shown in Figure 4.
Appendix B Implementation Details
B.1 Key Normalization
Transposition-equivariance is a crucial property in symbolic music: a composition’s musicality remains unchanged under global pitch shifts (uniform pitch adjustment across all notes). However, uneven key distribution in datasets can lead to data sparsity during training, potentially causing models to perform inconsistently across different keys and incorrectly assume key-dependent composition rules. Two approaches address this issue: (1) data augmentation through systematic semitone transpositions ${0,\pm 1,\pm 2,\cdots}$ Malandro [2024]; Zhao et al. [2024], or (2) normalizing all songs to a common key (e.g., C major and A minor) Lu et al. [2023]. We adopt the latter approach, implementing a modified version of Lu et al. [2023]’s method. Our implementation uses a key dictionary mapping 24 keys (12 major and 12 minor) to 12-dimensional binary vectors, where 1s indicate scale notes. Key detection is performed by computing dot products between these vectors and a song’s pitch histogram, with the highest-scoring key determining the transposition needed to normalize to C major or A minor. Further details can be found in the code.
B.2 Instrument Quantization
Band Arrangement
While our tokenization scheme supports all MIDI program IDs, many instruments with different IDs share fundamental compositional properties, differing primarily in timbre (e.g., acoustic and electric pianos). To leverage these similarities and reduce instrument distribution sparsity in training data, we group similar instruments and assign them the lowest program ID within their group. For multiple tracks of the same instrument type, we merge their notes into a single track.
Piano Arrangement
For piano arrangement, we consolidate all piano-type instruments (MIDI program IDs 0-7, including both acoustic and electric pianos) into a single track to form the target sequence.
B.3 Baseline Models
Tokenization Comparison
When comparing the effectiveness of REMI+ and the proposed REMI-z tokenization, the model structure used is the same as our arrangement model, but with a simpler training objective—the standard left-to-right language modeling.
Band Arrangement
For band arrangement, we use the Transformer-VAE model from Zhao et al. [2024] as our baseline, utilizing their official implementation without retraining. The model’s generation module was trained on Slakh2100 Manilow et al. [2019] (the same dataset as our fine-tuning) and its prior model leveraged the larger Lakh MIDI Dataset Raffel [2016], which encompasses Slakh2100.
Piano Reduction
Piano reductions can be categorized into two types: (1) piano accompaniments where the melody is delegated to a separate lead instrument, focusing solely on preserving harmony and texture (e.g., Zhao et al. [2023]; Wang et al. [2020]), which can be effectively handled using self-supervised methods; and (2) complete solo piano arrangements that additionally include the melody, requiring careful human arrangement and band-to-piano parallel datasets for supervised learning (e.g., Terao et al. [2023, 2022]). Due to the lack of open-source parallel datasets, we focus on the accompaniment arrangement task in this paper and use the term piano reduction interchangeably. However, our methodology is inherently flexible and can also handle the second type of reduction if parallel datasets become available. Additionally, exploring the potential for full solo piano reductions using unsupervised approaches remains a intriguing direction for future research.
For comparison, we reimplemented the UNet baseline from Terao et al. [2023] and trained it on Slakh2100 using our data preparation pipeline. Notably, we omitted octave shifts during input processing for two reasons: First, our output consistently represents a subset of the input notes, unlike their setup where human-composed piano references might use different octaves for the same notes. Second, introducing octave shifts in training significantly degraded performance, reducing note F1 from 58.31 to 42.80 and diminishing perceptual quality.
Drum Arrangement
For drum arrangement, we employ Composer’s Assistant v2 Malandro [2024] as our baseline, specifically using v2.1.0 from their official repository [^2] https://github.com/m-malandro/composers-assistant-REAPER. We use the model without retraining since its original training data (Lakh MIDI Dataset Raffel [2016]) encompasses our fine-tuning dataset (Slakh2100).
B.4 Hyperparameter Settings
Hyperparameters Values
seed 42
learning_rate 0.0005
weight_decay 0.1
train_batch_size 12
gradient_accumulation_steps 8
total_train_batch_size 96
optimizer
Adam with
betas=(0.9,0.999)
and epsilon=1e-08
lr_scheduler_type cosine
lr_scheduler_warmup_steps 1000
num_epochs 1
Table 7: Pre-train hyper-parameter setting.
The model we adopted is a GPT-2 model comprising 12 Transformer decoder layers with a hidden size of 768. We detail the hyperparameter configurations used in our experiments below.
For pre-training, Table 7 summarizes the pre-training configuration, where hyperparameters were used as-is without optimization. The pre-training was implemented using pytorch and transformers frameworks on a Linux platform, while fine-tuning additionally utilized lightning.
For fine-tuning, we conducted a simple learning rate search over 1e-5, 5e-5, 1e-4, selecting the optimal rate based on validation loss. This resulted in learning rates of 5e-5 for drum arrangement and 1e-4 for band arrangement and piano reduction. The batch sizes and context lengths were configured as follows: band arrangement used a batch size of 24 and context length of 768; piano reduction similarly adopted a batch size of 24 and context length of 768; drum arrangement employed a batch size of 8 and context length of 1536. Across all fine-tuning tasks, we used the AdamW optimizer with 0.01 weight decay, incorporating a linear learning rate scheduler with 500-step warmup. Training spanned 3 epochs for band, piano, and drum tasks, with early stopping patience of 2 epochs. The best checkpoints were selected based on validation loss performance.
For the tokenization comparison, the model is trained from scratch without generative pre-training. The hyperparameter settings remain the same as those used for band arrangement fine-tuning, except for an increased number of epochs (5).
All experiments were conducted with a fixed random seed of 42.
Appendix C Objective Metrics
In this section, we formally define the objective metrics used in the paper.
Most of our objective metrics are based on note-level F1 scores calculated on piano roll of 16-th note quantization. Let $X=\{x_{1},...,x_{n}\}$ and $Y=\{y_{1},...,y_{m}\}$ be two sequences of note events, where each note event $e$ is defined as a tuple $e=(t,p)$ with onset time $t$ and pitch $p$. A note event $x_{i}\in X$ is considered to match $y_{j}\in Y$ if and only if:
$|t_{x_{i}}-t_{y_{j}}|<\delta_{t}\text{ and }p_{x_{i}}=p_{y_{j}}$ (4)
where $\delta_{t}$ is the temporal tolerance threshold (set to one 16th-note duration in our evaluation).
Let $M(X,Y)$ denote the set of matched note pairs between $X$ and $Y$. The Note F1 score is defined as:
Precision $=\frac{|M(X,Y)|}{|X|}$ (5)
Recall $=\frac{|M(X,Y)|}{|Y|}$ (6)
Note F1 $=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (7)
The Notei F1 extends this by considering instrument matching, where each note event becomes $e=(t,p,i)$ with $i$ representing the instrument. The matching criterion becomes:
$|t_{x_{i}}-t_{y_{j}}|<\delta_{t}\text{ and }p_{x_{i}}=p_{y_{j}}\text{ and }i_{x_{i}}=i_{y_{j}}$ (8)
Notei F1 is then computed using Equation 7 with this stricter matching criterion.
For Melody F1 (Mel F1), given a multitrack piece with tracks $T=\{T_{1},...,T_{k}\}$, we first identify the melody track $T_{m}$ as:
$T_{m}=\operatorname*{arg\,max}_{T_{i}\in T}\frac{1}{|T_{i}|}\sum_{e\in T_{i}}p_{e}$ (9)
where $|T_{i}|$ is the number of notes in track $T_{i}$ and $p_{e}$ is the pitch of note event $e$. The Mel F1 score between output and target sequences is then:
[…]
For instrument control evaluation, Instrument IoU (I-IoU) measures the overlap between instrument sets. Let $\mathcal{I}^{out}$ and $\mathcal{I}^{tgt}$ denote the sets of instruments used in the output and target sequences respectively:
[…]
where an instrument is considered present if there exists at least one note event using it.
Finally, Voice WER (V-WER) evaluates the similarity of voice arrangements between two multitrack compositions. For each piece, we derive an ordered voice sequence by:
1) Computing the average pitch $\bar{p}_{i}$ for each active instrument $i\in\mathcal{I}^{active}$:
$\bar{p}_{i}=\frac{1}{|T_{i}|}\sum_{e\in T_{i}}p_{e}$ (12)
2) Constructing a voice sequence $V$ by sorting instruments by descending average pitch:
$V=[i_{1},i_{2},...,i_{n}]\text{ where }\bar{p}_{i_{k}}\geq\bar{p}_{i_{k+1}}\text{ for }k=1,...,n-1$ (13)
The V-WER between output sequence $V^{out}$ and target sequence $V^{tgt}$ is:
[…]
where $S$, $D$, and $I$ are the minimum number of substitutions, deletions, and insertions required to transform $V^{out}$ into $V^{tgt}$, and $N=|V^{tgt}|$ is the length of the target sequence.
Appendix D Subjective Evaluation
D.1 Subjective Metrics
These metrics are based on listeners’ auditory experiences and their subjective feelings while listening to the music. Since they are subjective, they cannot be easily defined using mathematical equations. However, we detail the evaluation criteria through the prompt questions provided in the questionnaire, which participants answered after listening to the demos. Each metric is assessed on a 5-point scale, ranging from 1 (very low) to 5 (very high).
  • Coherence: Does the arrangement flow naturally and smoothly? How consistent is each instrument’s performance and style throughout the piece?
  • Creativity: How creative is the arrangement while maintaining faithfulness and naturalness?
  • Musicality: What is the overall musical quality?
  • Faithfulness (band): How closely does the arrangement resemble the original piece in terms of melody and overall feel?
  • Faithfulness (piano): How closely does the arrangement capture the overall feel of the original piece?
  • Instrumentation (band-only): Does each instrument fulfill its appropriate role within the band, and do they harmonize effectively?
  • Playability (piano-only): How well is the piece suited for piano? How likely is it that a human pianist could perform this accompaniment?
  • Compatibility (drum-only): Is the drum beat compatible with the other instruments?
  • Phrase Transition (drum-only): How effectively does the drum arrangement handle transitions between phrases?
D.2 Human Evaluation Details
For the band arrangement task, we use an out-of-domain test with new compositions and new instrument groups for subjective evaluation, while for piano reduction and drum arrangement, we use songs from the test set to facilitate comparison with the original piano track (Rule-O) and the ground truth drum track. Each model/method was tasked with arranging the full song, and a phrase from the chorus, the most representative part, was selected for comparison, each containing 15s–30s of audio. We prepared 6 songs for the band arrangement, 5 for the piano reduction, and 5 for the drum arrangement. In the questionnaire, participants were first presented with the original song clip to be arranged, followed by the anonymized results from different models/methods in a random order. Participants were asked to evaluate each arranged clip on a 5-point scale, using the questions provided for each metric as outlined in §Subjective Metrics. The group of songs evaluated by each participant (containing one original arrangement and all corresponding model outputs) was also randomly selected. In total, we received 56, 73, and 77 groups of feedback for the band arrangement, piano reduction, and drum arrangement tasks, respectively.
Here are specific details for the band arrangement setting. In the band arrangement tasks, we tested the models’ ability in different scenarios by using various instrument settings. The content conditions were derived from either a piano arrangement or a full band arrangement. The instrument conditions included three different instrument sets, arranged in descending pitch order: string trio (violin, viola, cello), rock band (synth lead, clean electric guitar, distorted electric guitar, electric bass), and jazz band (saxophone, violin, brass section, clean electric guitar, piano, string ensemble, electric bass). Among the 6 songs used for the band arrangement task, two songs were arranged in each of the three settings. Additionally, the input came from songs composed for different settings, including 3 piano solos and 3 band arrangements.
Appendix E Probing Analysis of Pre-Training Impact
Model Acc@1 Acc@3 Acc@5
Random guess 2.94 8.82 14.71
Random initialized 38.98 61.71 74.22
FT only 41.50 64.93 76.53
PT only 46.14 69.30 79.61
PT + FT 45.89 68.96 79.47
Table 8: Instrument probing results.
Chord Root Chord quality
Model Acc@1 Acc@3 Acc@1 Acc@3
Random guess 8.33 25.00 11.11 33.33
Random initialized 48.86 78.76 38.05 77.97
FT only 50.09 80.23 39.26 79.03
PT only 62.93 89.42 50.09 85.58
PT + FT 58.05 85.93 44.92 82.48
Table 9: Chord probing results.
Knowledge probing techniques are used to explore what a model has learned within its hidden representations Rogers et al. [2020]. A widely adopted method is linear probing, where the pre-trained model is frozen and a simple linear classifier is trained on top of the hidden representations to predict specific properties, thereby revealing the extent to which particular knowledge is encoded in the model Alain and Bengio [2016]. This technique is particularly useful in our context for evaluating what the model has internalized during pre-training.
We have shown in the paper that the proposed models outperform the baseline models that do not undergo the pre-traning stage on the generation quality. In this section, we further analyze the reason. Specifically, we conduct probing experiments to assess whether pre-training enhances the acquisition of musical knowledge during fine-tuning to facilitate understanding content conditions in fine-tuning.
We focus on two probing tasks: (1) classifying instrument types from content sequences that contain only position, pitch, and duration tokens, and (2) recognizing chord progression sequences from content sequences. To determine whether instrument or chord information is linearly accessible within the model’s sequence embeddings, we employ linear classifier probes. In these probing experiments, the average pooling of the Transformer’s output embeddings across all tokens in a sequence is fed into the classifier. The model parameters are kept frozen, and only the linear classifiers are trained. We compare the knowledge captured by different models: a randomly initialized Transformer, a model that has undergone pre-training only (PT only), a model trained on the arrangement task without pre-training (FT only), and a model that has undergone both pre-training and fine-tuning (PT + FT).
For the probing experiments, the batch size was set to 12 for chord probing and 64 for instrument probing. The learning rates were 5e-4 for chord probing and 1e-4 for instrument probing. Both experiments shared the same remaining hyperparameters: training for 10 epochs, using a linear learning rate scheduler, a 500-step warmup, and a weight decay of 0.01.
E.1 Instrument Type Probing
In this task, we use a linear probe to estimate the instrument type from a single-track music sequence without providing instrument tokens. The goal is to predict which instrument is most likely to play the given note sequence. The performance is evaluated using top-1, top-3, and top-5 prediction accuracy metrics.
As shown in Table 8, a model initialized with random weights shows notable improvement after fine-tuning, suggesting that the ability to discern instrument styles is the requirement for performing well in music arrangement tasks. However, the gains in accuracy are modest, likely due to the limited number of training samples, which may constrain the model’s ability to aquire such knowledge through fine-tuning alone. Interestingly, models that underwent pre-training exhibit the highest accuracy in predicting instrument types. This indicates that substantial knowledge of instrument styles can be acquired effectively during the pre-training phase. Moreover, the proposed models that are both pre-trained and fine-tuned (PT+FT) maintained high accuracy levels, demonstrating that the knowledge about musical instruments is retained through the subsequent fine-tuning.
E.2 Chord Progression Probing
In this task, we use linear probes to predict chord progressions from a 2-bar music content sequence without instrument tokens. Eight linear probes are trained simultaneously to predict the chord roots and qualities for a total of four chords (two chords per bar). This tasks is used to evaluate whether the model contains position-specific chord information. Similarly, we use top-1 and top-3 prediction acuracy metrics.
As shown in Table 9, the model with only fine-tuning (FT only) indicates a foundational grasp of chord knowledge for this complex music analysis tasks. However, similar to instrument prediction, the knowledge of recognizing chord progression does not significant gain, until the pre-training is also introduced into the model, confirming that pre-training establishes a robust basis for better understanding of content sequence, which may potentially help with the quality of arrangement.